• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Calarco

    Log entry time 16:09:03 on January11,2002

    Entry number 80093

    keyword=Shift Summary

    The first commissioning shift for RCS ends with no target and no beam. The
    target crew is working on the software. At present, the prognosis of when it
    will work is unknown. Unfortunately, I am told that in order to continue the
    work on the software requires access to the hall which precludes working on
    beam delivery.

    The parasitic setup of HRS_r requires it to be set at such a low momentum
    that it cannot regulate within the required tolerance. Hence, the alarm is
    continuously on, an unacceptable condition for RCS since it would mask
    any alarms on the primary systems. John LeRose and Javier Gomez are
    working on the regulation system as I write this.

    We discovered that the scaler rates in S0 as seen in the HRS were higher
    (by approximately a factor of 2) than S0' as seen in the calorimeter. This
    was traced to the fact that the S0' discriminator threshold had never been
    set. It was at a default value of 100 mV. Lowering this to 30 mV doubled
    the scaler rate. Lowering it an additional 10 mV to a value of 20 mV only
    caused an increase of about 10%. Thus we conclude we are in the valley
    between the minimum ionizing peak and the PMT noise. So we will leave
    it here.

    Thus endeth the first day shift!